# This Time is Different: An Observability Perspective on Time Series Foundation Models

### v1

@@ 3.2 Training data
We trained Toto with a dataset of approximately 2.36 trillion time series points, of which 1.59 trillion are non-repeated and non-synthetic. This is significantly larger than the pretraining corpora of existing time series foundation models (Fig. 3A). Critically, 43% of our training mixture contains anonymous metrics from Datadog’s observability platform. This data excludes any customer data and is sourced solely from the our own monitoring of internal systems. It consists only of numerical time series data, and does not include any identifiable metadata. However, much of this data is sparse, noisy, or too granular or high in cardinality to be useful in its raw form. To curate a high-quality dataset, we sample queries based on quality and relevance signals from dashboards, monitor alerts, and notebooks built by domain experts using the platform.
Alongside the observability data, we include public time series datasets, in particular, the GIFT-Eval Pretrain [1] and Chronos [11] collections. Importantly, we remove the subset of the Chronos datasets that overlap with the GIFT-Eval benchmark in order to avoid any leakage from the test data. We also find that adding synthetic data improves model generalization and performance. For more details on the preparation of public, synthetic, and observability data, please see Appendix B.
@@ 4.1 Dataset
[…]
Like Toto’s pretraining data, the observability data comprising Boom is sourced from the Datadog platform. However, to ensure a robust evaluation setting and prevent contamination, training data for Toto (see Section 3.2) originates exclusively from production environments, while Boom’s evaluation data is drawn from a separate staging environment. Because isolation between production and staging data is guaranteed at the platform level, we can be certain there is no leakage. Like the training data, the Boom data excludes any customer data and is sourced solely from the platform’s own monitoring of internal systems.
Boom’s dataset consists of approximately 350 million data points across 2,807 metric queries generated through the monitoring of real-world distributed systems (Fig. 8). These series vary significantly in sampling frequency, temporal length, and dimensionality, capturing realistic operational conditions (Table 1). Despite having fewer unique series compared to GIFT-Eval (2,807 vs. 144K), Boom features substantially more total data points (350M vs. 158M) and significantly higher dimensionality, with a median of 60 variates per series compared to GIFT-Eval’s predominantly univariate or low-cardinality multivariate series. Both Boom and GIFT-Eval are significantly larger and more diverse than the older LSF benchmark. For accessibility in light of Boom’s large scale, we also provide (and evaluate on) a smaller uniformly sampled subset (“Boomlet”), which we describe in Appendix C.3.
@@ 4.3 Statistical characteristics
Observability time series are noted for having challenging distributional properties, including sparsity, spikes, and noisiness [70]; anomaly contamination [71] (which can manifest as skew and/or kurtosis); and for their high-cardinality multivariate nature, with complex relationships among variates [72]. They can also exhibit extreme dynamic range [31]; for example, metrics recording disk usage in bytes can range across many orders of magnitude. To illustrate these properties, we examine six relevant statistics for each variate in Boom, GIFT-Eval, and LSF (compared in Fig. 5):
@@
The ARCH-LM statistic [73] detects autoregressive conditional heteroskedasticity. Lower values suggest time-varying volatility; the Boom distribution is bimodal with a large peak near zero.
Spectral entropy [72], which quantifies the unpredictability of a series in the frequency domain, is also higher on average in observability series. This indicates less periodicity and greater irregularity.
The KPSS statistic [74], a test of nonstationarity, and assumes larger values in the Boom observability data, suggesting more frequent deviations from a deterministic trend.
“Flat spots” [75] measures the longest constant subsequence with a time series, normalized by series length. This metric is higher in Boom, indicative of sparse metrics within observability data.
@@ 4.4 Evaluation protocol
To evaluate models on Boom we closely follow the evaluation methodology proposed by Aksu et al. [1] for GIFT-Eval, including its standardized prediction lengths, strides, and train/validation/test splits. As in GIFT-Eval, our key accuracy metrics are Mean Absolute Scaled Error (MASE) [76] and approximate Continuous Ranked Probability Score (CRPS) [77, 78], both normalized by the Seasonal Naive forecast. And like GIFT-Eval, we also compute an average rank (where rank is computed as the mean rank with respect to CRPS across all prediction tasks).
We deviate from the GIFT-Eval protocol in two key respects in order to handle the presence of metrics with mostly constant values and infrequent spikes, which can cause problems in computing traditional metrics. First, we employ a shifted geometric mean for aggregating MASE and CRPS across tasks, which is stable in scenarios where constant subsequences would otherwise cause a standard geometric mean to collapse [79, 80]. Second, we use a separate evaluation approach for a small subset of the time series in Boom where extreme cases of constant subsequences still cause numerical instability for naive-forecast-normalized metrics. See Appendix C for details.
@@ 5 Experiments
Dataset # Series # Variates # Points Interval Series Length # Variates Per Series Pred. Length
@@
Boom 2,807 32,887 350 M 10 sec. 1 day 101 13,631 16,384 1 60 100 48 900
Boomlet 32 1,627 23M 10 sec. 1 hr. 5,231 16,384 16,384 21 49 100 48 900
GIFT-Eval 144,246 147,688 158 M 10 sec. 1 yr. 19 1,043 140,256 1 1 21 6 900
LSF 6 370 11 M 10 min. 1 hr. 17,420 39,500 69,680 7 7 321 96 720
Table 1: Key statistics of the three time series benchmarks used for evaluation in this study. Boom, introduced in this paper, contains approximately twice as many total time series points as GIFT-Eval (350M vs. 158M) and vastly more than LSF. Boom consists of multivariate series of much higher average cardinality than the other benchmarks, with a median of 60 variates per series vs. only 1 for GIFT-Eval and 7 for LSF. Conversely, Boom contains a narrower range of intervals compared to GIFT-Eval, ranging from 10 seconds to 1 day, vs. 10 seconds to 1 year for GIFT-Eval. This reflects the generally shorter time scales relevant to the observability domain.
@@
Boom. We evaluate Toto’s zero-shot forecasting performance alongside other foundation models, [11, 12, 13, 14, 23, 24], as well as full-shot statistical baselines. We do not evaluate full-shot deep learning models on Boom; as discussed in Section 1, these models are often impractical in real-word observability systems at scale. Similarly, we were unable to evaluate TabPFN [81] (a recent zero-shot model with strong performance on GIFT-Eval) on BOOM, as its open-source implementation’s lack of support for batched inference made it impractically slow to run. Details of the inference settings and evaluation procedures for all models are described in Appendix C.2. As shown in Table 2, Toto consistently outperforms other models, achieving 13.1% and 12.4% lower MASE and CRPS, respectively, than the next best (MoiraiBase), and a significantly lower rank (2.351 vs. 4.278). Similar qualitative results hold on the Boomlet subset (Appendix D.1.1).
@@
Dataset Metric Toto MoiraiBase TimesFM2.0 ChronosBolt-Base Time-MoELarge Timer VisionTS Auto-ARIMA Auto-ETS Auto-Theta
Boom MASE $\downarrow$ 0.617 0.710 $0.725$ $0.726$ $0.881$ $0.796$ $0.988$ $0.824$ $0.842$ $1.123$
CRPS $\downarrow$ 0.375 0.428 $0.447$ $0.451$ $0.643$ $0.639$ $0.673$ $0.736$ $1.975$ $1.018$
Rank $\downarrow$ 2.369 4.328 $5.243$ $5.576$ $9.843$ $9.920$ $10.989$ $9.686$ $11.671$ $12.472$
Boomlet MASE $\downarrow$ 0.617 0.779 0.685 $0.711$ $0.793$ $0.807$ $0.912$ $0.922$ $0.969$ $1.030$
CRPS $\downarrow$ 0.519 0.630 0.603 $0.637$ $0.780$ $0.793$ $0.885$ $0.880$ $15.664$ $1.182$
Rank $\downarrow$ 1.244 4.644 4.178 $5.578$ $9.500$ $9.844$ $11.722$ $10.278$ $12.856$ $13.289$
Table 2: Boom results. Performance of Toto, other zero-shot models, and baselines. MASE and CRPS are normalized by the Seasonal Naive forecast and aggregated across tasks using shifted geometric mean. Rank is the mean rank across tasks with respect to CRPS. For model families with multiple sizes (Moirai, Chronos) we show the variant that performs best on the full Boom evaluation. Toto significantly outperforms other methods on all metrics. We observe similar qualitative trends on the Boomlet results. Additional results, including all model sizes evaluated as well as categorical breakdowns, are available in Appendix D.1. Key: Best results, Second-best results.
@@
Table 3: GIFT-Eval results. Toto’s performance vs. the top Zero Shot, Full Shot, and Baseline models reported on the leaderboard [3]. MASE and CRPS are normalized by the Seasonal Naive forecast and aggregated across tasks using geometric mean. Rank is the mean rank across tasks with respect to CRPS. For multi-size models (Moirai, Chronos) we show the best-performing variant. Toto achieves top performance on GIFT-Eval in Rank and MASE, and second-best performance in CRPS. Key: Best results, Second-best results.
GIFT-Eval. We evaluate Toto’s zero-shot performance on general-purpose time series forecasting via the GIFT-Eval benchmark [1]. Toto achieves the top performance among all reported models, with an average ranking score of 5.495 as of May 2025. It achieves strong results both in point forecasting, with a MASE of 0.673, and probabilistic forecasting, with a CRPS of 0.437. Details of our evaluation settings for GIFT-Eval are provided in Appendix D.2. Notably, Toto is the top-performing method in spite of the fact that several competing models have known partial data leakage with the benchmark, as discussed by Aksu et al. [1].
Metric Toto MoiraiSmall MoiraiBase MoiraiLarge Time-MoEBase Time-MoELarge Time-MoEUltra VisionTS
MAE $\downarrow$ 0.324 0.357 0.344 0.350 - - - 0.345
MSE $\downarrow$ 0.303 0.327 0.310 0.330 - - - 0.309
Best Count 8 0 0 0 0 0 1 3
Table 4: LSF results Zero‐Shot comparison of models on the LSF benchmark. Non-Toto values are reproduced from published tables. Key: Best results, Second-best results. “-” used for models which were not evaluated on the Electricity dataset. “Best Count” row reports the number of times each model attains the best result for a given dataset-metric pair. Full results are available in Table 15.
LSF. We evaluate Toto on the widely-used Long Sequence Forecasting (LSF) benchmark [2]. Despite known limitations regarding dataset diversity and saturation by supervised methods [1, 29], Toto achieves state-of-the-art results in zero-shot evaluations, attaining the best performance on 8 out of 12 reported metrics when compared against other zero-shot methods, and the lowest average MAE and MSE, see Table 4. Furthermore, while our focus in this work is on the zero-shot setting, we explored the efficacy of fine-tuning Toto on the training splits of LSF and report the results in Table 16. We find that Toto achieves state-of-the-art results in full-shot evaluations, also attaining the best performance on 8 out of 12 reported metrics, and the lowest average MAE and MSE of all methods. These results underscore Toto’s strong generalization capabilities and its effectiveness when fine-tuned on specialized datasets, making it a versatile choice for a wide range of time series forecasting tasks. See Appendix D.3 for full results and additional experimental details.
@@ A.1 Input/output scaling
To ensure numerical stability, we compute the $\hat{\mu_{t}}$ and $\hat{s_{t}}$ using an efficient vectorized implementation (Listing 1) of Welford’s online algorithm [46], incorporating Bessel’s correction to provide an unbiased estimator of variance, as described in Option A of [82]. We stabilize training against extreme outliers by incorporating weak information from the global statistics.
@@ A.2 Attention mechanism
  • Concatenating variates along the time dimension and computing full self-attention between every variate/time location, as in the “any-variate attention” used by Woo et al. [13]. This can capture every possible variate and time interaction, but it is costly in terms of computation and memory usage.
  • Assuming variate independence, and computing attention only in the time dimension as in Nie et al. [19], Shi et al. [14]. This is efficient, but throws away all information about variate-wise interactions.
  • Computing attention only in the variate dimension, and using a feed-forward network in the time dimension [83, 35].
  • Computing “factorized attention,” where each transformer block contains a separate variate and time attention computation [47, 48, 49]. This allows both variate and time mixing, and is more efficient than full cross-attention. However, it doubles the effective depth of the network.

### v2

@@ 3.2 Training data
We trained Toto with a dataset of approximately 2.36 trillion time series points, of which 1.59 trillion are non-repeated and non-synthetic. This is significantly larger than the pretraining corpora of existing time series foundation models (Fig. 3A). Critically, 43% of our training mixture contains anonymous observability metrics from the Datadog platform. This data excludes any customer data and is sourced solely from the platform’s own monitoring of internal systems. It consists only of numerical time series data, and does not include any identifiable metadata. However, much of this data is sparse, noisy, or too granular or high in cardinality to be useful in its raw form. To curate a high-quality dataset, we sample queries based on quality and relevance signals from dashboards, monitor alerts, and notebooks built by domain experts using the platform.
Alongside the Datadog data, we include public time series datasets, in particular, the GIFT-Eval Pretrain [aksu2024giftevalbenchmarkgeneraltime] and Chronos [ansari2024chronoslearninglanguagetime] collections. Importantly, we remove the subset of the Chronos datasets that overlap with the GIFT-Eval benchmark in order to avoid any leakage from the test data. We also find that adding synthetic data improves model generalization and performance. For more details on the preparation of public, synthetic, and Datadog data, please see Appendix B.
@@ 4.1 Dataset
Like the Toto pretraining data, the observability data comprising Boom is also sourced from the Datadog platform. However, to ensure a robust evaluation setting and prevent contamination, training data for Toto (see Section 3.2) originates exclusively from production environments, while Boom’s evaluation data is drawn from a separate staging environment. Because isolation between production and staging data is guaranteed at the platform level, we can be certain there is no leakage. Like the training data, the Boom data excludes any customer data and is sourced solely from the platform’s own monitoring of internal systems.
Boom’s dataset consists of approximately 350 million data points across 2,807 metric queries generated through the monitoring of real-world distributed systems (Fig. 8). These series vary significantly in sampling frequency, temporal length, and cardinality, capturing realistic operational conditions (Table 1). Despite having fewer unique series compared to GIFT-Eval (2,807 vs. 144K), Boom features substantially more total data points (350M vs. 158M) and significantly higher dimensionality, with a median of 60 variates per series compared to GIFT-Eval’s predominantly univariate or low-cardinality multivariate series. Both Boom and GIFT-Eval are significantly larger and more diverse than the older LSF benchmark. For accessibility in light of Boom’s large scale, we also provide (and evaluate on) a smaller uniformly sampled subset (“Boomlet”), which we describe in Appendix C.3.
@@ 4.3 Statistical characteristics
Observability time series are noted for having challenging distributional properties, including sparsity, spikes, and noisiness [cheng2023aioperationsaiopscloud]; anomaly contamination [mou2025rocarobustcontrastiveoneclass] (which can manifest as skew and/or kurtosis); and for their high-cardinality multivariate nature, with complex relationships among variates [Palaskar_Ekambaram_Jati_Gantayat_Saha_Nagar_Nguyen_Dayama_Sindhgatta_Mohapatra_Kumar_Kalagnanam_Hemachandra_Rangaraj_2024]. They can also exhibit extreme dynamic range [Joosen_2023]; for example, metrics recording disk usage in bytes can range across many orders of magnitude. To illustrate these properties, we examine six relevant statistics for each variate in Boom, GIFT-Eval, and LSF (compared in Fig. 5):
@@
The ARCH-LM statistic [engle_autoregressive_1982] detects autoregressive conditional heteroskedasticity. Lower values suggest time-varying volatility; the Boom distribution is bimodal with a large peak near zero.
Spectral entropy [Palaskar_Ekambaram_Jati_Gantayat_Saha_Nagar_Nguyen_Dayama_Sindhgatta_Mohapatra_Kumar_Kalagnanam_Hemachandra_Rangaraj_2024], which quantifies the unpredictability of a series in the frequency domain, is also higher on average in observability series. This indicates less periodicity and greater irregularity.
The KPSS statistic [kwiatkowski_testing_1992], a test of nonstationarity, and assumes larger values in the Boom observability data, suggesting more frequent deviations from a deterministic trend.
“Flat spots” [tsfeatures_hyndman] measures the longest constant subsequence with a time series, normalized by series length. This metric is higher in Boom, indicative of sparse metrics within observability data.
@@ 4.4 Evaluation protocol
To evaluate models on Boom we closely follow the evaluation methodology proposed by aksu2024giftevalbenchmarkgeneraltime for GIFT-Eval, including its standardized prediction lengths, strides, and train/validation/test splits. As in GIFT-Eval, our key accuracy metrics are Mean Absolute Scaled Error (MASE) [Hyndman2006] and approximate Continuous Ranked Probability Score (CRPS) [GneitlingCRPS, parkCRPS], both normalized by the Seasonal Naive forecast. And like GIFT-Eval, we also compute an average rank (where rank is computed as the mean rank with respect to CRPS across all prediction tasks).
We deviate from the GIFT-Eval protocol in two key respects in order to handle the presence of metrics with mostly constant values and infrequent spikes, which can cause problems in computing traditional metrics. First, we employ a shifted geometric mean for aggregating MASE and CRPS across tasks, which is stable in scenarios where constant subsequences would otherwise cause a standard geometric mean to collapse [RavanaGmap, Beach2024]. Second, we use a separate evaluation approach for a small subset of the time series in Boom where extreme cases of constant subsequences still cause numerical instability for naive-forecast-normalized metrics. See Appendix C for details.
@@ 5 Experiments
Dataset # Series # Variates # Points (M) Interval Series Length # Variates Per Series Pred. Length
@@
Boom $2,807$ $32,887$ $350$ 10 sec. 1 day $101$ $13,631$ $16,384$ 11 $60$ $100$ $48$ $900$
Boomlet $32$ $1,627$ $23$ 10 sec. 1 hr. $5,231$ $16,384$ $16,384$ 2121 $49$ $100$ $48$ $900$
GIFT-Eval $144,246$ $147,688$ $158$ 10 sec. 1 yr. 1919 $1,043$ $140,256$ 11 11 2121 66 $900$
LSF 66 $370$ $11$ 10 min. 1 hr. $17,420$ $39,500$ $69,680$ 77 77 $321$ $96$ $720$
Table 1: Key statistics of the time series benchmarks used for evaluation in this study. Boom, introduced in this paper, contains approximately twice as many total time series points as GIFT-Eval (350M vs. 158M) and vastly more than LSF. Boom, consists of multivariate series of much higher average cardinality than the other benchmarks, with a median of 60 variates per series vs. only 1 for GIFT-Eval and 7 for LSF. Conversely, Boom contains a narrower range of intervals compared to GIFT-Eval, ranging from 10 seconds to 1 day, vs. 10 seconds to 1 year for GIFT-Eval. This reflects the generally shorter time scales relevant to the observability domain. Boomlet, a smaller subset of Boom that we provide for easier evaluation, exhibits similar statistics.
@@
Boom. We evaluate Toto’s zero-shot forecasting performance alongside other foundation models, [ansari2024chronoslearninglanguagetime, das2024a, Woo2024, shi2025timemoe, 10.5555/3692070.3693383, chen2024visionts], as well as full-shot statistical baselines. We do not evaluate full-shot deep learning models on Boom; as discussed in Section 1, these models are often impractical in real-word observability systems at scale. Similarly, we were unable to evaluate TabPFN [hoo2025tabpfn] (a recent zero-shot model with strong performance on GIFT-Eval) on BOOM, as its open-source implementation’s lack of support for batched inference made it impractically slow to run. Details of the inference settings and evaluation procedures for all models are described in Appendix C.2. As shown in Table 2, Toto consistently outperforms other models, achieving 13.1% and 12.4% lower MASE and CRPS, respectively, than the next best (MoiraiBase), and a significantly lower rank (2.351 vs. 4.278). Similar qualitative results hold on the Boomlet subset (Appendix D.1.1).
@@
Dataset Metric Toto MoiraiBase TimesFM2.0 ChronosBolt-Base Timer Time-MoEBase VisionTS Auto-ARIMA Auto-ETS Auto-Theta DLinear DeepAR
Boom MASE $\downarrow$ 0.617 0.710 0.725 0.726 0.796 0.881 0.988 0.824 0.842 1.123 - -
CRPS $\downarrow$ 0.375 0.428 0.447 0.451 0.639 0.643 0.673 0.736 1.975 1.018 - -
Rank $\downarrow$ 2.369 4.328 5.243 5.576 9.920 9.843 10.989 9.686 11.671 12.472 - -
Boomlet MASE $\downarrow$ 0.617 0.779 0.685 0.711 0.807 0.793 0.912 0.922 0.969 1.030 0.823 0.883
CRPS $\downarrow$ 0.519 0.630 0.603 0.637 0.793 0.780 0.885 0.880 15.664 1.182 0.641 0.697
Rank $\downarrow$ 1.300 5.244 4.844 6.411 11.344 11.011 13.278 11.778 14.156 14.522 6.044 7.822
Table 2: Boom results. Performance of Toto, other zero-shot models, and baselines. MASE and CRPS are normalized by the Seasonal Naive forecast and aggregated across tasks using shifted geometric mean. Rank is the mean rank across tasks with respect to CRPS. For model families with multiple sizes (Moirai, Chronos) we show the best-performing variant. Toto significantly outperforms other methods on all metrics. We observe similar qualitative trends on the Boomlet results. Additional results, including all model sizes evaluated as well as categorical breakdowns, are available in Appendix D.1. Key: Best results, Second-best results.
GIFT-Eval. We evaluate Toto’s zero-shot performance on general-purpose time series forecasting via the GIFT-Eval benchmark [aksu2024giftevalbenchmarkgeneraltime]. Toto achieves the top performance among all reported models, with an average ranking score of 5.495 as of May 2025. It achieves strong results both in point forecasting, with a MASE of 0.673, and probabilistic forecasting, with a CRPS of 0.437. Details of our evaluation settings for GIFT-Eval are provided in Appendix D.2. Notably, Toto is the top-performing method in spite of the fact that several competing models have known partial data leakage with the benchmark, as discussed by aksu2024giftevalbenchmarkgeneraltime.
@@
Table 3: GIFT-Eval results. Toto’s performance vs. the top Zero Shot, Full Shot, and Baseline models reported on the leaderboard [gifteval_leaderboard]. MASE and CRPS are normalized by the Seasonal Naive forecast and aggregated across tasks using geometric mean. Rank is the mean rank across tasks with respect to CRPS. For multi-size models (Moirai, Chronos) we show the best-performing variant. Toto achieves top performance on GIFT-Eval in Rank and MASE, and second-best performance in CRPS. Key: Best results, Second-best results.
LSF. We evaluate Toto on the widely-used Long Sequence Forecasting (LSF) benchmark [Wu2021]. Despite known limitations regarding dataset diversity and saturation by supervised methods [aksu2024giftevalbenchmarkgeneraltime, xu2024specialized], Toto achieves state-of-the-art results in zero-shot evaluations, attaining the best performance on 8 out of 12 reported metrics when compared against other zero-shot methods, and the lowest average MAE and MSE (see Table 4). Furthermore, while our focus in this work is on the zero-shot setting, we explored the efficacy of fine-tuning Toto on the training splits of LSF and report the results in Table 15. We find that Toto achieves state-of-the-art results in full-shot evaluations, also attaining the best performance on 8 out of 12 reported metrics, and the lowest average MAE and MSE of all methods. These results underscore Toto’s strong generalization capabilities and its effectiveness when fine-tuned on specialized datasets, making it a versatile choice for a wide range of time series forecasting tasks. See Appendix D.3 for full results and additional experimental details.
Zero Shot
Dataset Metric Toto MoiraiSmall MoiraiBase MoiraiLarge VisionTS Time-MoEBase Time-MoELarge Time-MoEUltra
ETTh1 MAE $\downarrow$ 0.413 0.424 0.438 0.469 0.414 0.424 0.419 0.426
MSE $\downarrow$ 0.435 0.400 0.434 0.510 0.390 0.400 0.394 0.412
ETTh2 MAE $\downarrow$ 0.363 0.379 0.382 0.376 0.375 0.404 0.415 0.399
MSE $\downarrow$ 0.340 0.341 0.345 0.354 0.333 0.366 0.405 0.371
ETTm1 MAE $\downarrow$ 0.378 0.409 0.388 0.389 0.372 0.415 0.405 0.391
MSE $\downarrow$ 0.396 0.448 0.381 0.390 0.374 0.394 0.376 0.356
ETTm2 MAE $\downarrow$ 0.303 0.341 0.321 0.320 0.321 0.365 0.361 0.344
MSE $\downarrow$ 0.267 0.300 0.272 0.276 0.282 0.317 0.316 0.288
Electricity MAE $\downarrow$ 0.243 0.320 0.274 0.273 0.294 - - -
MSE $\downarrow$ 0.161 0.233 0.188 0.188 0.207 - - -
Weather MAE $\downarrow$ 0.245 0.267 0.261 0.275 0.292 0.297 0.300 0.288
MSE $\downarrow$ 0.224 0.242 0.238 0.259 0.269 0.265 0.270 0.256
Mean MAE $\downarrow$ 0.324 0.357 0.344 0.350 0.345 - - -
MSE $\downarrow$ 0.304 0.327 0.310 0.330 0.309 - - -
Best Count 8 0 0 0 3 0 0 1
Table 4: LSF results Zero‐Shot comparison of models on the LSF benchmark. Non-Toto values are reproduced from published tables. Key: Best results, Second-best results. “Best Count” row reports the number of times each model attains the best result for a given dataset-metric pair.
@@ A.1 Input/output scaling
To ensure numerical stability, we compute the $\hat{\mu_{t}}$ and $\hat{s_{t}}$ using an efficient vectorized implementation (Listing 1) of Welford’s online algorithm [welford1962note], incorporating Bessel’s correction to provide an unbiased estimator of variance, as described in Option A of [stackexchange_variance]. We stabilize training against extreme outliers by incorporating weak information from the global statistics.
@@ A.2 Attention mechanism
  • Concatenating variates along the time dimension and computing full self-attention between every variate/time location, as in the “any-variate attention” used by Woo2024. This can capture every possible variate and time interaction, but it is costly in terms of computation and memory usage.
  • Assuming variate independence, and computing attention only in the time dimension as in Nie2023, shi2025timemoe. This is efficient, but throws away all information about variate-wise interactions.
  • Computing attention only in the variate dimension, and using a feed-forward network in the time dimension [ilbert2024samformer, Liu2024].
  • Computing “factorized attention,” where each transformer block contains a separate variate and time attention computation [zhang2023crossformer, Rao2021, Arnab2021]. This allows both variate and time mixing, and is more efficient than full cross-attention. However, it doubles the effective depth of the network.